OpenStudy (anonymous):

If P(x) = 4x3 - 5x2 + 8x + 7, find P(-3).

OpenStudy (saifoo.khan):

Substitute -3 instead of x.

OpenStudy (callisto):

P(-3) = 4(-3)^3 - 5(-3)^2 + 8(-3) + 7 = -108 - 45 -24 +7 = -170